OpenStudy (anonymous):

Which measure is the same as the standard error of the mean? standard deviation of the population standard deviation of the sample means mean of the sample mean of the population

jimthompson5910 (jim_thompson5910):

what are your thoughts on this?

OpenStudy (anonymous):

i have no clue

jimthompson5910 (jim_thompson5910):

The standard error is another name for the standard deviation of the sample mean (xbar) distribution

jimthompson5910 (jim_thompson5910):

Standard Error (SE): \[\Large SE = \frac{\sigma}{\sqrt{n}}\] this is when sigma is known/given

jimthompson5910 (jim_thompson5910):

if sigma is unknown, but s is known, then \[\Large SE = \frac{s}{\sqrt{n}}\]